Fluffy the lemming is wandering the northern tundra when two angry moose kick him like a soccer ball. The first moose kicks Fluffy with a velocity of 20.0 m/s [W] and the second moose kicks him back at 16.0 m/s [E]. What is Fluff’s acceleration if it takes him 0.005 min to travel between the two moose?
1
Expert's answer
2020-10-01T10:10:06-0400

First, convert 0.005 minutes to seconds: 0.005*60=0.3 s.

Next, remember that acceleration is how quickly velocity changes. We have 20 m/s West and then 16 m/s East (in opposite direction), which means that the change in velocity is 36 m/s:


"a=\frac{v_\text{final}-v_\text{initial}}{t}=\frac{36}{0.3}=120\text{ m/s}^2."

It was a miracle if Fluffy survived this.